# FAST-GSC: Fast and Adaptive Semantic Transmission for Generative Semantic Communication

### v1

@@ III-A Semantic Extraction
We denote the source image as ${\mathbf{s}}\in{\mathbb{R}^{c\times h\times w}}$. The semantic extraction process of distilling the $k$th semantic unit ${{\mathbf{u}}_{k}}$ can be expressed by
${{\mathbf{u}}_{k}}={f_{k}}\left({{\mathbf{s}};{{\bm{\psi}}_{k}}}\right),$ (1)
@@
where ${{{\bm{\psi}}_{k}}}$ represents the trainable parameters of the extraction model for distilling ${{\mathbf{u}}_{k}}$, $K$ represents the maximum number of semantic units and $\forall k\in\left\{{0,1,...,K-1}\right\}$.
III-B Diffusion-based Inference
Inspired by the significant success of diffusion models across a wide range of real-world generative inference tasks, the Stable Diffusion model [18] is commonly employed for generative inference at the receiver within the GSC framework [2, 3]. Its goal is to generate images that semantically adhere to the provided prompt.
[…]
During sampling, the noise estimation can be performed by classifier-free guidance [29], which is expressed by
where $w$ is the guidance scale.
As DDPM necessitates executing all consecutive denoising steps to produce the clean sample ${{{\mathbf{x}}_{0}}}$, it results in significant task latency. In light of this, we adopt the sampling method of the denoising diffusion implicit model (DDIM) [30], which allows for skipping some denoising steps without compromising the quality of the inferred images. This method enhances DDPM by leveraging a non-Markovian inference process, which can be formulated as
${{\mathbf{x}}_{t-1}}=$ $\sqrt{{{\bar{\alpha}}_{t-1}}}\left({\frac{{{{\mathbf{x}}_{t}}-\sqrt{1-{{\bar{\alpha}}_{t}}}{{\tilde{\varepsilon}}_{\theta}}\left({{{\mathbf{x}}_{t}},{\mathbf{y}},t}\right)}}{{\sqrt{{{\bar{\alpha}}_{t}}}}}}\right)$
$+\sqrt{1-{{\bar{\alpha}}_{t-1}}}\cdot{{\tilde{\varepsilon}}_{\theta}}\left({{{\mathbf{x}}_{t}},{\mathbf{y}},t}\right).$ (5)
III-C Latency Discussion and Communication Mechanism
The task latency can be categorized into the following three components:
1) Semantic Extraction Latency: Due to limited computational resources, the extraction process proceeds sequentially. Consequently, the aggregate latency of the entire extraction procedure can be expressed as
${\tau_{ext}}=\sum\limits_{k=0}^{K-1}{{\tau_{k}}},$ (6)
where ${{\tau_{k}}}$ is the latency for extracting the $k$th semantic unit. In practice, ${\tau_{k}}$ hinges on the efficiency of each extraction model and the computational power available at the transmitter, with durations extending from milliseconds to seconds [31, 32, 33].
2) Diffusion-based Inference Latency: In the inference process, the latency is determined by the total number of denoising steps $M$, which can be formulated as
${\tau_{gen}}=\sum\limits_{m=0}^{M-1}{{\tau_{m}}},$ (7)
where ${{\tau_{m}}}$ is the inference latency of the denoising step $m$ and $m\in\left\{{0,1,...,M-1}\right\}$. In practice, ${\tau_{m}}$ depends on the computational resources of the receiver, varying from milliseconds to seconds [34, 35].
3) Wireless Transmission Latency: Generally, the prompts distilled in GSC systems are concise and highly compressed [2, 3]. Consequently, leveraging the capabilities of 5G technology can significantly reduce the transmission latency, ${\tau_{trans}}$, to less than one millisecond [36].
Based on the investigations above, it is evident that the latencies associated with extraction and inference significantly exceed those of transmission. Therefore, the system latency is primarily influenced by the extraction and inference processes. Under complex requirements, extensive semantic unit extraction is essential to thoroughly capture the semantic attributes of the source. Concurrently, executing a high number of denoising steps is crucial to ensure the precision of inference results. Consequently, the combination of these procedures leads to significant system latency.
In conventional GSC, the aforementioned processes operate in a temporally disjointed manner, resulting in cumulative latency. By contrast, in PGSC, we propose executing the semantic extraction and diffusion-based inference in parallel. Specifically, the diffusion model’s denoising process is initiated by the first received semantic units. In subsequent stages, newly arrived semantic units combine with previously arrived units to collectively guide the denoising process. A comparison of the communication mechanisms and system latency of the conventional GSC and our PGSC is presented in Fig. 1.
Figure 1: Comparison of communication mechanisms and system latency. (a) Conventional GSC framework, where transmission starts after the completion of all semantic extraction; (b) Proposed PGSC framework, where semantic extraction and diffusion-based inference are executed in parallel.
IV From PGSC to FAST-GSC: Motivations and Approaches
In this section, we commence by outlining our refinement motivations, which are based on our examination of the PGSC framework’s implementation. Subsequently, we elucidate our enhancements from the perspectives of both the transmitter and the receiver.
IV-A Refinement Motivations
Figure 2: Illustrative examples of some direct implementations of the PGSC framework. The total denoising step is set at 30 for all schemes. In (a), the switch of the execution order of these two extraction models yields semantically different results. In (b), the switch of the execution order of the two extraction models doesn’t change the semantic meanings of inference results. In (c), we posit that the speed of the emotion recognition model on the upper side is slower than on the lower side, leading to the emotion-related semantic unit on the upper side arriving 10 steps later than on the lower side. Observations indicate that the inferred image on the lower side effectively incorporates guidance from both extracted semantic units, whereas the upper one fails to integrate the second semantic feature.
Although the PGSC framework allows for continuous conditioning on received semantic units, thereby reducing system latency resulting from the prolonged extraction process, this approach can potentially lead to significant task performance loss. To better illustrate, we present three representative examples illustrated in Fig. 2 and summarize two key factors determining the task performance of PGSC as follows:
@@ IV-A Refinement Motivations
1) Sequential Dependencies Among Various Units: Given that different semantic units influence varying aspects and extents of the inference results, they also exhibit diverse degrees of sequential dependencies during the inference process. As illustrated in Fig. 2 (a), with the same extraction speed, changing the extraction and transmission sequence between two semantic units can yield significantly different outcomes. This phenomenon demonstrates the tight sequential dependency between these two semantic units during the inference process of the diffusion model. In contrast, as shown in Fig. 2 (b), switching the extraction and transmission sequence between another pair of semantic units results in similar inference outcomes both semantically and visually, indicating a weak sequential dependency between these units.
2) Reception Immediacy of Individual Units: In the PGSC framework, the denoising process begins after the reception of the first semantic unit. As the denoising process progresses, the noise level diminishes, limiting the capacity to incorporate new semantic features. As depicted in the upper part of Fig. 2 (c), due to the excessive processing delays of the second extraction model, the receiver can only integrate the second semantic unit into the conditional guidance at the 20th denoising step. The minimal remaining noise at the 20th denoising step limits the modification capacity of the second semantic unit, resulting in a certain degree of semantic mismatch between the source and inferred images. If the reception timestamp of the second semantic unit is 10 denoising steps earlier, its influence can still be observed, as shown in the lower part of Fig. 2 (c).
In light of these analyses, we propose enhancing the system performance of the PGSC framework from two perspectives: 1) From the transmitter’s perspective, we aim to learn the intrinsic sequential dependencies among various semantic units and strategically design their extraction and transmission order accordingly; and 2) From the receiver’s perspective, we aim to mitigate the stringent reception immediacy and enhance the guidance effectiveness of the late-arrived semantic units for the final performance. The details of these two approaches are presented in the following content.
IV-B Perspective from Transmitter: Temporal Prompt Engineering
@@
At the transmitter, our goal is to guide the diffusion model to produce the desired output by regulating the input. This is akin to the objective of prompt engineering [22], except that we need to achieve this from a temporal perspective, forming a temporal prompt engineering approach. Given our observation that the order of extraction and transmission of different semantic units should follow their sequential dependencies, we resort to RL to learn and implement this strategy. Its aim is to ensure the final inferred images bear a close resemblance to the comprehensive prompt, while simultaneously minimizing system latency as much as possible. The designs of the fundamental components are outlined in the following content.
@@
1) MDP Constructing: In constructing the RL model, we frame the interaction between the agent and the environment as a Markov Decision Process (MDP). The MDP can be defined as a three-component tuple $\left({\mathcal{S},\mathcal{A},r}\right)$, where $\mathcal{S}$ is state space set, $\mathcal{A}$ is the action space set, and $r$ is the reward function. Within this framework, an episode is defined as a complete trajectory from the initial state to the terminal state, consisting of a sequence of states, actions, and corresponding rewards. To simplify the learning process of RL models, we divide the inference process into multiple segments at equal intervals. At the beginning of each inference segment, only newly arrived semantic units can be integrated with previously arrived ones to guide the denoising. This approach allows the RL model to learn reception deadlines of semantic units with different attributes, thereby establishing the extraction and transmission sequence. To better illustrate the MDP in our study, we depict an episode in Fig. 3 and the detailed design of this framework is given as follows:
@@
State: We include the task-related request in the agent’s state, which contains the attribute and quantity of semantic units required by the receiver. To facilitate the learning of the RL model, we can employ one-hot encoding to preprocess the semantic attributes. Let $\mathbf{e}_{k,t}$ denote the encoded $N_{\text{e}}$-length one-hot vector for the $k$-th required semantic attribute at phase [^1] In our work, the term ‘phase’ corresponds to a single iteration or transition within the RL environment, with each phase forming a segment of an episode. $t$, where the $n_{\text{e}}$-th position is set to 1, corresponding to the activation of the $n_{\text{e}}$-th extraction model to obtain the relevant semantic unit. When the requested number of semantic units is less than $K$, we fill the remaining positions in the one-hot encoding with zero vectors. Subsequently, the entire encoded one-hot matrix for the whole prompt can be expressed by $\mathbf{E}_{t}=\left[\mathbf{e}_{0,t},\ldots,\mathbf{e}_{K-1,t}\right]$. Since the residual noise levels vary at different denoising steps, and the allowable space for modifications also changes, the decision-making process of the RL model should include the conditional denoising step of the diffusion model $d_{t}$ as part of its state, which indicates the starting denoising step from which the semantic units selected at time $t$ can guide. When $d_{t}$ reaches the designated maximum number of denoising steps, it signifies the conclusion of the episode. To distinguish whether the semantic units are still required by the receiver at phase $t$, we also design a binary $K$-length vector ${{{\mathbf{c}}_{t}}}$ as an indicator. When an element of this vector is set to 0, it indicates that the corresponding semantic unit has either been transmitted or the associated one-hot attribute vector is a zero vector. The element is set to 1 otherwise. Therefore, the corresponding state observed by the agent at time phase $t$ is denoted as ${{\mathbf{s}}_{t}}=\left[{{{\mathbf{E}}_{t}},{{\mathbf{c}}_{t}},{d_{t}}}\right]$.
@@
Action: The action executed by the agent at time $t$ is denoted by ${{\mathbf{a}}_{t}}=\left[{{a_{0}},...,{a_{K-1}}}\right]$, where ${a_{k}}\in\left\{{0,1}\right\}$ for $k=0,...,K-1$. The practical implication of this action is that, when the agent selects certain semantic attributes, the corresponding extraction models that distill the semantic units with the selected attributes are executed at this RL phase.
@@
Reward: Aligned with our objectives of ensuring high inference precision while reducing task latency, the reward is a composite of two components: the system’s performance reward and the latency reward. Similar to [2], we adopt the CLIP score [37] as the performance reward in our work. This metric assesses the semantic similarity between the inferred images and the corresponding prompts within the embedding space of CLIP model [38], which can be expressed by
@@
Building on the PGSC framework, system latency is determined by the degree of parallel execution between semantic extraction and inference processes. Specifically, in the first phase of each episode, the receiver has not yet received any semantic units, and thus, denoising has not yet started. The latency of this phase is entirely determined by the execution time of the selected extraction model. From the second phase onward, the receiver can condition on the received semantic units, and extraction at the transmitter and inference at the receiver start simultaneously. The latency of these phases is determined by the maximum latency of either the transmitter’s semantic extraction or the receiver’s segmented denoising. Thus, instead of calculating system latency at the conclusion of one episode, we can expedite the learning process by defining the latency reward as the additional extraction delay relative to the segmented denoising duration at each phase. Therefore, the combined reward obtained at phase $t$ can be formulated as
@@
To achieve this, we utilize the invalid action masking technique [42], which is commonly implemented to avoid repeatedly generating invalid actions in large discrete action spaces. Specifically, we set the logits associated with the aforementioned non-contributory actions to $\infty$- ∞ and normalize the probabilities of all the actions by
@@
IV-C Perspective from Receiver: Sequential Conditional Denoising
@@
Figure 4: Demonstration of the refinements at the receiver. (a) Workflow of the proposed semantic difference calculation module; (b) Illustration of the sequential conditional denoising process, where the guidance effectiveness of late-arrived semantic units is enhanced for improved final performance.
@@
At the receiver, we aim to mitigate the stringent reception immediacy, thereby enabling the sequentially arrived units to continuously contribute to the inference results. To achieve this goal, one straightforward method is to add a certain amount of noise to the inference process every time a new semantic unit arrives at the receiver, ensuring enough space for modifications. However, this approach increases the overall noise level, requiring additional denoising steps and significantly prolonging the inference process.
@@
Inspired by the classifier-free guidance [29], we propose a sequential conditioning denoising approach to enhance the intervention capability of late-arrived semantic units for final performance without disrupting the ongoing denoising schedule. Specifically, we first introduce a semantic difference calculation module that can output the difference between the estimated conditional noises guided by the newly received and the previously received semantic features, as depicted in Fig. 4 (a). Contrary to blindly introducing random noise to all the intermediate results, this module allows us to preciously and efficiently identify the space the newly arrived semantic features target to modify at the current denoising step. Subsequently, we supplement the noise difference to the estimated noise conditioned on all the received semantic features using a weighted sum method. Building upon Equation (4), this modification can be formulated as follows:
@@
where $\alpha$ is the intervention factor, ${{{\mathbf{y}}_{{\text{combined}}}}}$ denotes all the semantic units received at diffusion step $t$, ${{{\mathbf{y}}_{{\text{new}}}}}$ denotes the semantic units newly received at diffusion step $t$ and ${{{\mathbf{y}}_{{\text{previous}}}}}$ denotes the semantic units received before diffusion step $t$. Through this approach, the guidance effectiveness of late-arrived semantic units for final performance can be enhanced, which can mitigate the issue of many semantic features not being effectively incorporated into the inference results due to their late extraction and transmission in complex tasks. The overall sequential conditional denoising process is illustrated in Fig. 4 (b).

### v2

@@ IV-A Temporal Prompt Engineering
We denote the source collected from sensor $i$-th as ${\mathbf{s}_{i}}\in{\mathbb{R}^{c\times h\times w}}$. The semantic extraction process of employing the $k$-th extraction model for distilling the semantic units ${\mathbf{u}}_{ki}$} from $\mathbf{s}_{i}$} can be expressed by
${{\mathbf{u}}_{ki}}={f_{k}}\left({{\mathbf{s}_{i}};{{\bm{\psi}}_{k}}}\right),$ (1)
@@
where ${\bm{\psi}}_{k}$}} represents the trained parameters of the extraction model $k$, $K$ represents the maximum number of the semantic extraction models and $\forall k\in\left\{{0,1,...,K-1}\right\}$.
However, the execution of a large number of extraction models imposes significant challenges in computation resource-constrained scenarios, potentially resulting in prolonged transmission latency and excessive computational memory consumption. Therefore, we propose to execute the semantic extraction models sequentially to distill key semantic units for transmission.
Rather than collecting all semantic units for every user prior to transmission, the semantic units distilled at the current timestamp undergo a multi-user resource allocation procedure and are transmitted to their respective target users immediately. Meanwhile, the diffusion model’s denoising process is initiated by the first received semantic units. In subsequent timestamps, newly arrived semantic units combine with previously arrived units to collectively guide the denoising process.
Figure 2: Illustrative examples of some direct implementations of the sequential transmission method. The total denoising step is set at 30 for all schemes. In (a), the switch of the execution order of these two extraction models yields semantically different results. In (b), the switch of the execution order of the two extraction models doesn’t change the semantic meanings of reconstruction results.
@@
However, initial experimental results indicate the presence of sequential dependencies among semantic units, suggesting that the execution order of extraction models significantly affects the quality of the reconstructed data. As illustrated in Fig. 2 (a), with the same extraction speed, changing the extraction and transmission sequence between two semantic units can yield significantly different outcomes. This phenomenon demonstrates the tight sequential dependency between these two semantic units during the reconstruction process of the diffusion model. In contrast, as shown in Fig. 2 (b), switching the extraction and transmission sequence between another pair of semantic units results in similar reconstruction outcomes both semantically and visually, indicating a weak sequential dependency between these units.
@@
Given the aforementioned observation that the order of extraction and transmission of different semantic units should follow their sequential dependencies, we resort to RL to learn and implement this strategy. Its aim is to ensure the final inferred images bear a close resemblance to the comprehensive prompt, while simultaneously minimizing system latency as much as possible. The designs of the fundamental components are outlined in the following content.
@@
1) MDP Constructing: In constructing the RL model, we frame the interaction between the agent and the environment as a Markov Decision Process (MDP). The MDP can be defined as a three-component tuple $\left({\mathcal{S},\mathcal{A},r}\right)$, where $\mathcal{S}$ is state space set, $\mathcal{A}$ is the action space set, and $r$ is the reward function. Within this framework, an episode is defined as a complete trajectory from the initial state to the terminal state, consisting of a sequence of states, actions, and corresponding rewards. To simplify the learning process of RL models, we divide the reconstruction process into multiple segments at equal intervals. At the beginning of each reconstruction segment, only newly arrived semantic units can be integrated with previously arrived ones to guide the denoising. This approach allows the RL model to learn reception deadlines of semantic units with different attributes, thereby establishing the extraction and transmission sequence. To better illustrate the MDP in our study, we depict an episode in Fig. 3 and the detailed design of this framework is given as follows:
@@
State: We include the task-related request in the agent’s state, which contains the attribute and quantity of semantic attributes required by the receiver. To facilitate the learning of the RL model, we can employ one-hot encoding to preprocess the semantic attributes. Let $\mathbf{e}_{k,t}$ denote the encoded $N_{\text{e}}$-length one-hot vector for the $k$-th required semantic attribute at phase [^1] In our work, the term ‘phase’ corresponds to a single iteration or transition within the RL environment, with each phase forming a segment of an episode. $t$, where the $n_{\text{e}}$-th position is set to 1, corresponding to the activation of the $n_{\text{e}}$-th extraction model to obtain the relevant semantic unit. When the requested number of semantic units is less than $K$, we fill the remaining positions in the one-hot encoding with zero vectors. Subsequently, the entire encoded one-hot matrix for the whole prompt can be expressed by $\mathbf{E}_{t}=\left[\mathbf{e}_{0,t},\ldots,\mathbf{e}_{K-1,t}\right]$. Since the residual noise levels vary at different denoising steps, and the allowable space for modifications also changes, the decision-making process of the RL model should include the conditional denoising step of the diffusion model $d_{t}$ as part of its state, which indicates the starting denoising step from which the semantic units selected at time $t$ can guide. When $d_{t}$ reaches the designated maximum number of denoising steps, it signifies the conclusion of the episode. To distinguish whether the semantic units are still required by the receiver at phase $t$, we also design a binary $K$-length vector ${\mathbf{c}}_{t}$}} as an indicator. When an element of this vector is set to 0, it indicates that the corresponding semantic unit has either been transmitted or the associated one-hot attribute vector is a zero vector. The element is set to 1 otherwise. Therefore, the corresponding state observed by the agent at time phase $t$ is denoted as ${{\mathbf{s}}_{t}}=\left[{{{\mathbf{E}}_{t}},{{\mathbf{c}}_{t}},{d_{t}}}\right]$.
@@
Action: The action executed by the agent at time $t$ is denoted by ${{\mathbf{a}}_{t}}=\left[{{a_{0}},...,{a_{K-1}}}\right]$, where ${a_{k}}\in\left\{{0,1}\right\}$ for k=0,,K1k=0,...,K-1. The practical implication of this action is that, when the agent selects certain semantic attributes, the corresponding extraction models that distill the semantic units with the selected attributes are executed at this RL phase.
@@
Reward: Aligned with our objectives of ensuring high reconstruction precision while reducing task latency, the reward is a composite of two components: the system’s performance reward and the latency reward. Similar to [2], we adopt the CLIP score [46] as the performance reward in our work. This metric assesses the semantic similarity between the inferred images and the corresponding prompts within the embedding space of CLIP model [47], which can be expressed by
@@
Building on this framework, system latency is determined by the degree of parallel execution between semantic extraction and reconstruction processes. Specifically, in the first phase of each episode, the receiver has not yet received any semantic units, and thus, denoising has not yet started. The latency of this phase is entirely determined by the execution time of the selected extraction model. From the second phase onward, the receiver can condition on the received semantic units, and extraction at the transmitter and reconstruction at the receiver start simultaneously. The latency of these phases is determined by the maximum latency of either the transmitter’s semantic extraction or the receiver’s segmented denoising. Thus, instead of calculating system latency at the conclusion of one episode, we can expedite the learning process by defining the latency reward as the additional extraction delay relative to the segmented denoising duration at each phase. Therefore, the combined reward obtained at phase $t$ can be formulated as
@@
To achieve this, we utilize the invalid action masking technique [51], which is commonly implemented to avoid repeatedly generating invalid actions in large discrete action spaces. Specifically, we set the logits associated with the aforementioned non-contributory actions to -$\infty$ and normalize the probabilities of all the actions by
@@
IV-B Sequential Conditional Denoising
Inspired by the significant success of diffusion models across a wide range of real-world generative reconstruction tasks, the Stable Diffusion model [31] is commonly employed for generative reconstruction at the receiver within the GSC framework [2, 3]. Its goal is to generate images that semantically adhere to the provided prompt.
The stable diffusion model is built upon the denoising diffusion probabilistic model (DDPM) [18], which is capable of learning a distribution ${p_{\theta}}\left({{{\mathbf{x}}_{0}}}\right)$ that approximates $q\left({{{\mathbf{x}}_{0}}}\right)$. It can be treated as a latent variable model of the form ${p_{\theta}}\left({{{\mathbf{x}}_{0}}}\right)=\int{{p_{\theta}}\left({{{\mathbf{x}}_{0:T}}}\right)}d{{\mathbf{x}}_{1:T}}$, where ${{\mathbf{x}}_{1}},...,{{\mathbf{x}}_{T}}$ are called latents of the same dimensionality as the original data ${{\mathbf{x}}_{0}}\sim q\left({{{\mathbf{x}}_{0}}}\right)$. The joint distribution ${p_{\theta}}\left({{{\mathbf{x}}_{0:T}}}\right)$ is called the reverse process, which can be formulated as ${p_{\theta}}\left({{{\mathbf{x}}_{0:T}}}\right)=p\left({{{\mathbf{x}}_{T}}}\right)\mathop{\prod}\limits_{t=1}^{T}{p_{\theta}}\left({\left.{{{\mathbf{x}}_{t-1}}}\right|{{\mathbf{x}}_{t}}}\right)$.
Conventionally, during the sampling process, noise estimation is typically performed using classifier-free guidance [52], which can be expressed as
where $w$ is the guidance scale.
Figure 4: Illustrative examples of some direct implementations of the traditional denoising method. The total denoising step is set at 30 for all schemes. We posit that the speed of the emotion recognition model on the upper side is slower than on the lower side, leading to the emotion-related semantic unit on the upper side arriving 10 steps later than on the lower side. Observations indicate that the inferred image on the lower side effectively incorporates guidance from both extracted semantic units, whereas the upper one fails to integrate the second semantic feature.
In FAST, the denoising process begins after the reception of the first semantic unit. As the denoising process progresses, the noise level diminishes, limiting the capacity to incorporate new semantic features. We conducted experiments to evaluate whether the traditional denoising method mentioned above is compatible with our transmission schemes discussed in Section IV-A. As depicted in the upper part of Fig. 4, due to the excessive processing delays of the second extraction model, the receiver can only integrate the second semantic unit into the conditional guidance at the 20-th denoising step. The minimal remaining noise at the 20-th denoising step limits the modification capacity of the second semantic unit, resulting in a certain degree of semantic mismatch between the source and inferred images. If the reception timestamp of the second semantic unit is 10 denoising steps earlier, its influence can still be observed, as shown in the lower part of Fig. 4. The results indicate that the reception immediacy of semantic units significantly impacts the quality of the reconstructed data.
@@ IV-B Sequential Conditional Denoising
Figure 5: Demonstration of the data reconstruction process. (a) Workflow of the proposed semantic difference calculation subunit; (b) Illustration of the sequential conditional denoising module, where the guidance effectiveness of late-arrived semantic units is enhanced for improved final performance.
@@
Therefore, we aim to mitigate the stringent reception immediacy at each user end, thereby enabling the sequentially arrived units to continuously contribute to the reconstruction results. To achieve this goal, one straightforward method is to add a certain amount of noise to the reconstruction process every time a new semantic unit arrives at the receiver, ensuring enough space for modifications. However, this approach increases the overall noise level, requiring additional denoising steps and significantly prolonging the reconstruction process.
@@
In contrast, we devise a sequential conditional denoising module to enhance the intervention capability of late-arrived semantic units for final performance without disrupting the ongoing denoising schedule. Specifically, we first introduce a semantic difference calculation subunit that can output the difference between the estimated conditional noises guided by the newly received and the previously received semantic features, as depicted in Fig. 5 (a). Contrary to blindly introducing random noise to all the intermediate results, this subunit allows us to preciously and efficiently identify the space that the newly arrived semantic units target to modify at the current denoising step. Subsequently, we supplement the noise difference to the estimated noise conditioned on all the received semantic units using a weighted sum method. Building upon classifier-free guidance [52] method, this modification can be formulated as follows:
@@
where $\alpha$ is the intervention factor, ${\mathbf{y}}_{{\text{combined}}}$}} denotes all the semantic units received at diffusion step $t$, ${\mathbf{y}}_{{\text{new}}}$}} denotes the semantic units newly received at diffusion step $t$ and ${\mathbf{y}}_{{\text{previous}}}$}} denotes the semantic units received before diffusion step $t$.
Through this approach, the guidance effectiveness of late-arrived semantic units for final performance can be enhanced, which can mitigate the issue of many semantic features not being effectively incorporated into the reconstruction results due to their late extraction and transmission in complex tasks. The sequential conditional denoising module and the overall data reconstruction process are illustrated in Fig. 5 (b).
IV-C Semantic-aware Resource Allocation
In conventional multi-user transmission systems, resource allocation is predominantly based on channel state information for bandwidth management. While effective for maximizing physical-layer metrics, they are insufficient for multi-user semantic transmission. This is because they fail to incorporate two critical dimensions of the problem: the complex semantic dependencies among information units, which dictate the value and decoding order of data, and the inherent disparity in task importance across users, which demands differentiated transmission priorities. Therefore, a novel strategy that jointly optimizes resources based on channel quality, semantic dependencies, and task priorities is essential.
Herein, we design an RL-based semantic-aware resource allocation module for dispatching the extracted semantic units. The key components are defined as follows:
State: For each user, we first compute the volume of the remaining extracted semantic units in the current time slot, categorized by semantic attribute, and store these values in a vector. We then append the user’s signal-to-noise ratio to the vector, followed by an indicator reflecting the user’s task importance. The state representation is formed by concatenating the vectors from all users. Assuming there are $K$ distinctive semantic attributes and $M$ users, the resulting state vector has a dimensionality of M(K+2)M*(K+2).
Action: The action executed by the agent at time $t$ is denoted as ${{\mathbf{a}}_{t}}=\left[{{a_{0}},...,{a_{M-1}}}\right]$, where $a_{m}$} stands for normalized bandwidth allocation and ${a_{m}}\in\left[{0,1}\right]$ for m=0,,M1m=0,...,M-1.
Reward: Aligned with the object of ensuring higher task performance for users with greater task importance, the initial reward is defined as the weighted sum of all users’ task performance:
${r_{s}}=\frac{1}{M}\sum\limits_{m=0}^{M-1}{{w_{m}}}*{{r_{p,m}}},$ (12)
where $r_{p,m}$}} denotes the performance score of user $m$, and $w_{m}$}} represents the task importance of user-$m$. Same as the performance reward defined in the temporal prompt engineering module, $r_{p,m}$}} can only be obtained at the end of the episode.
Informed by the temporal prompt engineering module, we devise a reward shaping strategy by measuring the semantic dependency of each semantic unit. Semantic units that have a higher probability of being extracted first are defined as possessing the highest degree of semantic dependence. The intermediate reward is defined as the weighted sum of transmitted semantic dependence at this step, which is shown as follows:
${r_{d}}=\frac{1}{M}\sum\limits_{m=0}^{M-1}{\sum\limits_{k=0}^{K-1}{{{w_{m}}}*{n_{m,k}}*{d_{m,k}}}},$ (13)
where $n_{m,k}$}} is the number of user-$m$’s transmitted units of attribute $k$ and $d_{m,k}$}} is the semantic dependence of attribute $k$.
Therefore, the combined reward obtained at step $t$ can be formulated as
[…]
where $L$ denotes the number of steps in one episode.
Building upon the MDP framework outlined above, we employ the PPO method described in Section IV-A to determine the optimized resource allocation strategy. A detailed discussion of this approach is omitted here for brevity.